OpenStudy (anonymous):

Point A is located (-2,2) and D is located at (4,2). Find the coordinates of the point that lies halfway between A and D. a) (-1,1) b) (0,1) c) (1,0) d) (1,1) medal and fan will be given

OpenStudy (anonymous):

@Hero hi can you help me with this please

OpenStudy (anonymous):

@jdoe0001 hey can you help me with this please

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

how did u get that ?

OpenStudy (anonymous):

@Hero any ideas?

OpenStudy (anonymous):

ok thanks

OpenStudy (haseeb96):

do u know how too chooose the correct answer

hero (hero):

@Haseeb96, I just warned you about giving answers.

OpenStudy (haseeb96):

okay

OpenStudy (anonymous):

yea and explaination would be great lol

OpenStudy (phi):

are you sure there is not a typo in the question or the answer choices? the correct answer is not listed. the x coordinate of the mid point is the average of the x values , the y coordinate of the mid-point is the average of the y values

hero (hero):

The midpoint formula is \(M = \dfrac{x_1 + x_2}{2}, \dfrac{y_1 + y_2}{2}\)

OpenStudy (anonymous):

there is no typo but its okay thanks for the help

hero (hero):

@gabs15, there's a reason @phi mentioned something about a typo.

OpenStudy (anonymous):

no... i didnt mean it in a way that he is wrong i mean that there is no typo that i wrote it exactly how the problem was @Hero @phi

OpenStudy (phi):

I assume you found the average of the x values, -2 and 4 which is (-2+4)/2 = 2/2 = 1 and the average of the y values 2 and 2 = (2+2)/2 = 4/2 = 2 (though we should know that the average of the same number is just that number, and do not need to use the formula) at any rate, the mid-point is (1,2) So mention this to your teacher. take a screen shot and show them the problem

OpenStudy (anonymous):

i will show her thanks